Within 11 years, the age of Peter will be half the square of the age he was 13 years ago. Calcute the current age of age
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

Within 11 years, the age of Peter will be half the square of the age he was 13 years ago.
Calculate the current age of Peter
:
a + 11 = .5(a-13)^2
a + 11 = .5(a^2 - 26a + 169)
multiply both sides by 2
2a + 22 = a^2 - 26a + 169
arrange as a quadratic equation
0 = a^2 - 26a - 2a + 169 - 22
a^2 - 28a + 147 = 0
Factors to
(a-7)(a-21) = 0
The reasonable solution
a = 21 yr is Peter's age
:
:
:
See if that works
a + 11 = .5(21-13)^2
a + 11 = .5(8^2)
a + 11 = .5(64)
a + 11 = 32
Yes, 21 + 11 = 32

Answer by ikleyn(52778)

.

For your info :


        "within"  means  "inside of something",  or  "between".

        This word is  INAPPLICABLE  in the context of your problem.

        You should write  "In 11 years . . . ",  which means  "11  years later",  or  "after  11  years".